# EVER: Exact Volumetric Ellipsoid Rendering for Real-time View Synthesis

### v5

@@ Abstract
We present Exact Volumetric Ellipsoid Rendering (EVER), a method for real-time differentiable emission-only volume rendering. Unlike recent rasterization based approach by 3D Gaussian Splatting (3DGS), our primitive based representation allows for exact volume rendering, rather than alpha compositing 3D Gaussian billboards. As such, unlike 3DGS our formulation does not suffer from popping artifacts and view dependent density, but still achieves frame rates of $\sim\!30$ FPS at 720p on an NVIDIA RTX4090. Since our approach is built upon ray tracing it enables effects such as defocus blur and camera distortion (e.g. such as from fisheye cameras), which are difficult to achieve by rasterization. We show that our method is more accurate with fewer blending issues than 3DGS and follow-up work on view-consistent rendering, especially on the challenging large-scale scenes from the Zip-NeRF dataset where it achieves sharpest results among real-time techniques.
Figure 1: An overview of the quality benefits of our EVER technique. Left: On the Zip-NeRF dataset, our model produces sharper and more accurate renderings than 3DGS and successor splatting-based techniques. Middle: Because our method correctly blends primitive colors according to the physics of volume rendering, it produces fewer “foggy” artifacts than splatting. Right: Our method correctly blends primitive colors, which is not possible using a splatting regardless of how primitives are sorted (globally or ray-wise).
@@ 1 Introduction
Unlike NeRF, 3DGS lacks a true volumetric density field, and uses Gaussians to describe the opacity of the scene rather than of density. As such, 3DGS’s scene representation may violate volumetric rendering, in that an anisotropic Gaussian primitive in 3DGS will appear to have the same opacity regardless of the viewing direction of the camera. This lack of a consistent underlying density field prohibits the use of various algorithms and regularizers (such as the distortion loss from Mip-NeRF360 [2]), but the biggest downside of this model is popping. Popping in 3DGS is due to two assumptions made by the model: that primitives do not overlap, and that (given a camera position) primitives can be sorted accurately using only their centers. These assumptions are almost always violated in practice, which causes the rendered image to change significantly as the camera moves due to the sort-order of primitives changing. This popping may not be noticeable when primitives are small, but describing a large scene using many small primitives requires a prohibitively large amount of memory.
In this work we build on the primitive based representation of 3DGS, but we define a physically accurate, constant density ellipsoid based representation. This formulation lets us compute the exact form of the volume rendering integral efficiently, which eliminates the inconsistencies of 3DGS while still maintaining real-time framerates. Because our system is built around ray tracing (rather than 3DGS’s rasterization) it can straightforwardly model various optical effects like radial distortion lenses including fisheye and defocus blur. Our method guarantees 3D consistent real-time rendering while also improving the image quality of our 3DGS baselines. This is particularly pronounced in the challenging large-scale Zip-NeRF scenes [4] where our method matches the quality of state-of-the-art offline rendering methods.
@@ 2 Related Work
Refer to caption Refer to caption Refer to caption Refer to caption
@@ 3 Motivation
The parameterization of the density and color fields can vary from small MLPs [29] to large hierarchies of hashes and grids grids [31]. Though this approach can produce highly realistic renderings, accurately integrating Equation 1 requires a large number of calls to the underlying field, which means that training and rendering may be slow.
@@
Constant Density Ellipsoids    Our model lies in between NeRF and 3DGS: like NeRF we perform 3D consistent volume rendering of a radiance field, but like 3DGS we parameterize the scene using a collection of anisotropic primitives. Instead of Gaussians as primitives we use constant density ellipsoids, which we can volume render exactly without any numerical quadrature (and efficiently thanks to decades of progress in ray tracing). To highlight the difference between our method, 3DGS, StopThePop, and traditional NeRF, we render the same 3 colorful ellipsoids with all 4 methods on the right side of Figure 1. When the primitives are sorted per a view, the color between the primitives do not mix to produce purple, yellow, and blue green, but the appearance is smooth. When the primitives are sorted per a ray, the order changes throughout the image, causing lines to appear. Because our rendering model is exact, the rendering from our model exactly matches the ground truth.
@@ 4 Method
Like 3DGS, our method takes as input a set of posed images and a sparse point cloud. We optimize a mixture of ellipsoids (each with a constant density and color) to reproduce the appearance of the input images, where the initial positions of ellipsoids are determined by the input point cloud. We build upon the 3DGS framework, and reuse its Adaptive Density Control (ADC), with some modifications for handling density based primitives. We begin by describing how we are able to perform exact volume rendering of our representation, then describe how we convert opacity to density to allow the optimization of our density based primitives.
@@ 4.1 Exact Primitive-based Rendering
Figure 3: A visualization of our rendering procedure. The volume rendering equation can be integrated a field of constant density ellipsoids (depicted at the top) by tracing 9 rays. When the ray enters each primitive, the density along the ray increases. When it exits, the density drops back down a corresponding amount.
@@
We use a simple primitive-based rendering model, where each primitive has a constant density and a constant (view-dependent) color. We choose our primitive’s shape to be an ellipsoid, which, similar to the Gaussians in 3DGS, is fully characterized by a rotation and scale matrix. We illustrate these primitives and their intersections in Figure 3.
Selecting density and color models that are piecewise-constant in 3D means that the density and color along any ray are also piecewise constant, and they can be written in the following step function form with coefficients $\sigma_{i},c_{i}$ respectively:
$\sigma(t)$ $=\sum_{i=1}^{N}\sigma_{i}\mathbbm{1}[t_{i}\leq t<t_{i+1}],$ (2)
$\mathbf{c}(t)$ $=\sum_{i=1}^{N}\mathbf{c}_{i}\mathbbm{1}[t_{i}\leq t<t_{i+1}],$ (3)
where $\mathbbm{1}[\cdot]$ is an indicator function whose value is $1$ when the condition in the brackets is met and 0 otherwise, and $\{t_{i}\}_{i=1}^{N}$ are the density discontinuities, which are found by intersecting the ray with the primitives. The density and color coefficients for the $i$th interval, $t\in[t_{i},t_{i+1})$, can be determined by combining the contributions of each primitive along the ray:
$\sigma_{i}=\sum_{k=1}^{i}\Delta\sigma_{k},\quad\mathbf{c}_{i}=\frac{1}{\sigma_{i}}\sum_{k=1}^{i}\Delta\sigma_{k}\Delta\mathbf{c}_{k},$ (4)
where $\Delta\sigma_{k}$ and $\Delta\mathbf{c}_{k}$ are the changes in density and color at the $k$th intersection point of the ray with the primitives in the scene, respectively. Note that $\Delta\sigma_{k}$ is positive when entering a primitive and negative when exiting one. Now, the volume rendering integral from Equation 1 can be computed exactly:
$C=\sum_{i=1}^{N}\mathbf{c}_{i}(1-\exp(-\sigma_{i}\Delta t_{i}))\prod_{j=1}^{i-1}\exp(-\sigma_{j}\Delta t_{j})$ (5)
@@
Unsurprisingly, the expression in Equation 5 is identical to the quadrature rule derived by Max [28] that also makes a piecewise-constant approximation to density and color — though in our case this piecewise-constant property is an intentional design decision, not an approximation.
Though our scene representation looks superficially similar to that of 3DGS, it is different in two key ways: First, while 3DGS Gaussians are treated as 2D “billboards”, our ellipsoids interact with each other so as to constitute a proper and consistent 3D radiance field. Second, while the 2D opacity profile of the primitives used in 3DGS always has a smooth Gaussian falloff, the 2D opacity profile of one of our ellipsoidal primitives can range from extremely smooth to a perfect step function in the limit of infinite density. See Figure 4 for examples of our primitive’s opacity profiles for different density values. Our ellipsoidal primitives are also beneficial due to their computational efficiency in intersection calculations, low memory footprint, and smooth surface geometry.
@@
4.2 Density Parameterization
As described in Section 4.1, our method assigns a density value to each primitive. Optimizing the density of all primitives directly presents a challenge: as the density of a primitive grows and its opacity approaches $1$, the gradient used for updating the parameters of the primitive approaches 0 and would require special treatment [25]. Intuitively, this is because the outside of the primitive becomes opaque.
We avoid this problem by optimizing over a parameter $\alpha$ and using the following density during rendering:
$\sigma(\alpha)=-\frac{\log(1-0.99\cdot\alpha)}{\min(s_{x},s_{y},s_{z})}\,.$ (6)
With this, a primitive with density $\sigma(\alpha)$ when viewed along it’s shortest axis will have a peak opacity equal to $\alpha$.
@@ 5 Implementation
Our model was implemented using PyTorch, CUDA, OptiX [33], and Slang [20]. We use OptiX to ray trace through the primitives and to provide a per-ray sorting of distances. During training we rebuild our BVH at every step, which takes tens of milliseconds. All shaders are written in Slang [20], a language that supports automatic differentiation with the Slang.D extension [1]. We use adjoint rendering to propagate the gradient, which means we skip storage of all function values and simply recompute them during the backwards pass. To optimize the representation, we use our differentiable renderer in the 3DGS [22] codebase, and make a few adjustments to handle density based primitives (Section 5.2 and Appendix C).
@@ 5.1 Ray Tracing (BVHs) for Sorting
We use ray tracing to sort primitives since it offers more flexibility than rasterization. Ray tracing allows for effects like fisheye projection and defocus blur (see Figure 5) as well as random pixel offsets during training, which we find improve performance.
We use the recent approach of 3DGRT [30], which uses a BVH to accelerate an exact per ray sort of the primitives to realtime speeds. We start by providing the Nvidia OptiX [33] framework with an Axis-Aligned Bounding Box (AABB) and an intersection function that indicates where along the ray it intersects each ellipsoid. A tight AABB for an ellipsoid is straightforward to compute, and we interesect a ray with an anisotropic ellipsoid first making it isotropic through shifting, rotating, then scaling the ray origin. Then, then we compute the stable isotropic ray/sphere intersection described Haines et al [19].
@@
where $\alpha\in[0,1]$ is opacity and $\mathbf{s}\in\mathbb{R}^{3}$ is axis-aligned scaled for each primitive. This regularizer is applied only to primitives that are visible in the current batch.
@@ 5.2 Changes to Adaptive Density Control
The Adaptive Density Control (ADC) used by 3DGS is critical to its success. During training 3DGS uses a variety of strategies for periodically cloning, splitting, and pruning 3D Gaussians. This is essential to avoid local minima by creating primitives where they are necessary, and removing primitives where they are invisible. Though these heuristics were developed specifically for 3DGS, they work well for our method with some minor necessary adjustments.

### v6

@@ Abstract
We present Exact Volumetric Ellipsoid Rendering (EVER), a method for real-time 3D reconstruction. EVER accurately blends an unlimited number of overlapping primitives together in 3D space, eliminating the popping artifacts that 3D Gaussian Splatting (3DGS) and other related methods exhibit. EVER represents a radiance field as a set of constant-density volumetric ellipsoids, which are raytraced by intersecting each primitive twice (once upon ray entrance and another on ray exit) and accumulating the derivatives of the densities and colors along the ray. Because EVER is built around ray tracing, it also enables effects such as defocus blur and fish-eye camera distortion, while still achieving frame rates of $\sim\!30$ FPS at 720p on an NVIDIA RTX4090. We show that our method is more accurate on the challenging large-scale scenes from the Zip-NeRF dataset, where it achieves state of the art SSIM, even higher than Zip-NeRF.
Figure 1: An overview of the quality benefits of our EVER technique. Left: On the Zip-NeRF dataset, our model produces the sharpest results ever. Middle: Our method is able to blend primitive colors together to reproduce light and shadow better than other 3DGS-based methods. Right: The way our approach blends colors is physically accurate over any number of primitives, matching the ground truth computed via quadrature.
@@ 1 Introduction
Unlike NeRF, 3DGS lacks a true volumetric density field, and uses Gaussians to describe the opacity of the scene rather than of density. As such, 3DGS’s scene representation isn’t volumetric rendering, as an anisotropic Gaussian primitive in 3DGS will appear to have the same opacity regardless of the viewing direction of the camera. This lack of a consistent underlying density field prohibits the use of various algorithms and regularizers (such as the distortion loss from Mip-NeRF360 [2]), but the biggest downside of this model is popping. Popping in 3DGS is due to two assumptions made by the model: that primitives do not overlap, and that (given a camera position) primitives can be sorted accurately using only their centers. These assumptions are almost always violated in practice, which causes the rendered image to change significantly as the camera moves due to the sort-order of primitives changing. This popping may not be noticeable when primitives are small, but describing a large scene using many small primitives requires a prohibitively large amount of memory.
In this work we build on the primitive based representation of 3DGS, but introduce a method that allows for the physically accurate blending of an infinite number of overlapping number of primitives, specifically constant density ellipsoids. We implement this blending method in a ray-tracing framework, which allows us to model various optical effects like radial distortion lenses including fisheye and defocus blur in a straightforward way, all at real-time framerates. Our method guarantees 3D consistent real-time rendering while also improving the image quality of our 3DGS baselines. This is particularly pronounced in the challenging large-scale Zip-NeRF scenes [4] where our method matches the quality of state-of-the-art offline rendering methods.
@@ 2 Related Work
Refer to caption $\bar{\downarrow}$ Refer to caption Refer to caption Refer to caption
@@ 3 Motivation
The parameterization of the density and color fields can vary from small MLPs [31] to large hierarchies of hashes and grids [33]. Though this approach can produce highly realistic renderings, accurately integrating Equation 1 requires a large number of calls to the underlying field, which means that training and rendering may be slow.
@@ 4 Method
Like 3DGS, our method takes as input a set of posed images and a sparse point cloud. We optimize a mixture of ellipsoids (each with a constant density and color) to reproduce the appearance of the input images. Our method enables exact volume rendering. In Section 4.1 we describe our scene representation and the algorithm for volume rendering it, and in Section 4.2 we describe how we optimize it to the input images.
@@ 4.1 Exact Primitive-based Rendering
Figure 3: A visualization of our blending algorithm. 1. We start with our primitives, in this case two, and we are trying to recover the value of the intervals along the ray. 2. We take the change in density and color and sort then from closest to furthest. 3. We sum these changes as we progress through the ray, retrieving the interval values.
@@
We use a simple primitive-based rendering model, where each primitive has a constant density and a single color (in 3D space). We choose our primitive’s shape to be an ellipsoid, which, similar to the Gaussians in 3DGS, is fully characterized by a position vector, rotation quaternion, and scale vector. To model view-dependent color, the ellipsoid’s color changes with the view direction, which we store as a spherical function represented by the coefficients for the first two degrees of spherical harmonics.
To render a given ray, we iterate over all front and back surfaces of all primitives intersected by the ray. When we step through the front of a primitive, we add its density $\Delta\sigma_{k}$ and premultiplied color $\Delta\sigma\Delta\mathbf{c}_{k}$ to two running totals, $\sigma_{i},\hat{\mathbf{c}}_{i}$. When we step through the back, we subtract the same density and pre-multiplied color. Using these running totals, we can calculate the color $\mathbf{c}_{i}$ and density $\sigma_{i}$ for the $i$th interval:
$\sigma_{i}=\sum_{k=1}^{i}\Delta\sigma_{k}\,,\quad\mathbf{c}_{i}=\frac{1}{\sigma_{i}}\hat{\mathbf{c}}_{i}=\frac{1}{\sigma_{i}}\sum_{k=1}^{i}\Delta\sigma_{k}\Delta\mathbf{c}_{k}\,.$ (2)
The key here is that this decoupling of the primitive into the front and back allow us to rearrange the surfaces along the ray, as seen in Figure 3. In this example, we can see two overlapping primitives. In 3DGS, we would simply render A and B separately, ignoring their overlap, and then combine them. In contrast, our method allows rendering this representation exactly by computing the density and color at the three intervals A, A+B, and B, by tracking their changes as we progress through the ray. Since color and density are constant within each interval of length $\Delta t_{i}$ between consecutive intersection points, there is a simple closed-form solution to the volume rendering equation:
$\mathbf{c}_{i}(1-\exp(-\sigma_{i}\Delta t_{i}))\,$, (3)
By tracking the running total and the distance between the intersection steps, $\Delta t_{i}$, we are able to integrate the volume rendering equation exactly by alpha compositing these simple segments, as follows:
$C=\sum_{i=1}^{N}\mathbf{c}_{i}(1-\exp(-\sigma_{i}\Delta t_{i}))\prod_{j=1}^{i-1}\exp(-\sigma_{j}\Delta t_{j})\,.$ (4)
@@
Unsurprisingly, the expression in Equation 4 is identical to the quadrature rule derived by Max [30], which uses a piecewise-constant approximation to density and color — though in our case this piecewise-constant property is an intentional design decision, not an approximation. Our work uses this closed-form solution to enable real-time rendering of the entire collection of ellipsoids without introducing any approximation errors which may cause artifacts.
@@
Note that although our scene representation looks superficially similar to that of 3DGS, it is different in two key ways: First, while 3DGS Gaussians are treated as 2D “billboards”, our ellipsoids blend together so as to constitute a proper and consistent 3D radiance field. Second, while the 2D opacity profile of the primitives used in 3DGS always has a smooth Gaussian falloff, the 2D opacity profile of one of our ellipsoidal primitives can range from extremely smooth to a perfect step function in the limit of infinite density. See Figure 4 for examples of our primitive’s opacity profiles for different density values.
4.2 Optimization
Our optimization framework is built on top of 3DGS, where instead of rasterizing Gaussians we ray trace and volume render our ellipsoid-based representation using the method described in Section 4.1. In particular, we reuse the Adaptive Density Control (ADC) introduced in 3DGS, with some modifications for our density-based primitives.
The first, essential change is to convert from the 3DGS opacity framework to a density framework. The naive approach would be to assign each primitive a density instead of opacity, but this presents a challenge. As the density of a primitive grows and its opacity approaches $1$, the gradient used for updating the parameters of the primitive approaches 0 and would require special treatment [27]. Intuitively, this is because the outside of the primitive becomes opaque.
We avoid this problem by optimizing over a parameter $\alpha_{k}$ which is mapped to density value used during rendering:
$\sigma(\alpha_{k})=-\frac{\log(1-0.99\cdot\alpha_{k})}{\min(s_{kx},s_{ky},s_{kz})}\,.$ (5)
With this, a primitive with density $\sigma(\alpha_{k})$ when viewed along its shortest axis will have a peak opacity equal to $\alpha_{k}$, ensuring at least one view will have non-zero gradient.
Finally, we add an additional splitting condition: We split primitives if the opacity is near 1 across the entire primitive when viewed from the major axis to see if the gradient has vanished using the following test:
$0.99<1-\exp\mathopen{}\left(-\sigma(\alpha_{k})\cdot\max\mathopen{}\left(s_{kx},s_{ky},s_{kz}\mathclose{}\right)\mathclose{}\right)\,.$ (6)
@@ 5 Implementation
Our model was implemented using PyTorch, CUDA, OptiX [35], and Slang [20]. We use OptiX to ray trace through the primitives and to provide a per-ray sorting of distances. During training we rebuild our BVH at every step, which takes tens of milliseconds. All shaders are written in Slang [20], a language that supports automatic differentiation with the Slang.D extension [1]. We use adjoint rendering to propagate the gradient, which means we skip storage of all function values and simply recompute them during the backwards pass. To optimize the representation, we use our differentiable renderer in the 3DGS [23] codebase, and make a few adjustments to handle density based primitives (Section 5.2 and Appendix B).
@@ 5.1 Ray Tracing (BVHs) for Sorting
We use ray tracing to sort primitives since it offers more flexibility than a rasterization approach like StP [36]. Ray tracing allows for effects like fisheye projection and defocus blur (see Figure 5) as well as random pixel offsets during training, which we find improves performance.
We start by providing the NVIDIA OptiX [35] framework with an Axis-Aligned Bounding Box (AABB) for the ellipsoid, which is used to construct a binary tree. We then provide an intersection test to refine the search if a ray hits the bounding box, for which we adapt the stable isotropic ray/sphere intersection described by Haines et al [19]. To further accelerate rendering, we also integrate the recent approach of 3DGRT [32], which accumulates multiple hits within a queue during each call to OptiX trace.
@@
where $\alpha\in[0,1]$ is the opacity and $\mathbf{s}\in\mathbb{R}^{3}$ is the ellipsoid axis scaling vector for each primitive. This regularizer is applied only to primitives that are visible in the current batch.
@@ 5.2 Changes to Adaptive Density Control
The Adaptive Density Control (ADC) used by 3DGS is critical to its success. During training 3DGS uses a variety of strategies for periodically cloning, splitting, and pruning 3D Gaussians. This is essential to avoid local minima by creating primitives where they are necessary and removing primitives where they are invisible. Though these heuristics were developed specifically for 3DGS, they work well for our method with some minor necessary adjustments.
To decide where to create primitives, 3DGS accumulates the $xy$ gradients of each primitive, then splits or clones them if they are above a certain threshold every $n$ steps. The spatial gradients of our model tend to be smaller, so the thresholds used by these splitting and cloning heuristics must be modified (splitting is $2.5\times 10^{-7}$, cloning is $0.1$). We use a similar splitting method as 3DGS: the primitive size is reduced and the new position is perturbed by some random amount sampled from a normal distribution, but we additionally also divide the density in half, similar to [7].
6 Results
Mip-NeRF360 Zip-NeRF Tanks&Temples & DeepBlending
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3DGS [23] 27.48 .816 .216 25.84 .817 .358 26.65 .848 .263
StopThePop [36] 27.33 .816 .212 25.92 .819 .352 26.60 .847 .252
3DGRT [32] 27.20 .818 .248 - - - 26.22 .865 .254
SMERF [15] 27.99 .818 .238 27.28 .829 .339 - - -
Our model 27.51 .825 .194 26.58 .845 .308 26.59 .889 .234
ZipNeRF [4] 28.54 .828 .198 27.37 .836 .305 - - -
GPU-hr $\downarrow$ Mem. $\downarrow$ FPS $\uparrow$ GPU-hr $\downarrow$ Mem. $\downarrow$ FPS $\uparrow$ GPU-hr $\downarrow$ Mem. $\downarrow$ FPS $\uparrow$
3DGS [23] 0.54 763MB 224 1.07  [^2] MB 559 0.34 563MB 560
StopThePop [36] 0.60 780MB 180 0.24 223MB 403 0.39 549MB 179
3DGRT [32] 0.83 383MB 156 - - - 0.83 388MB 190
SMERF [15] 272 139MB 454 528 4108MB 408 - - -
Our model 1.04 1134MB 36 1.26 1694MB 24 0.86 1173MB 44
ZipNeRF [4] 32 903MB  0.5 48 903M  0.5 - - -
Table 1: Results on the 9 scenes from Mip-NeRF360 [2], 4 large scenes from Zip-NeRF [4] datasets, and 4 scenes from DeepBlending [21] and Tanks&Temples [25] combined. The increased sharpness and perceptual quality are reflected in the qualitative comparison in Fig. 6. “GPU-hr” indicates the number of GPU hours it takes to train a model. “Mem.” indicates the amount of memory the finished model consumes on disk. Variations in memory usage between 3DGS, StopThePop, 3DGRT, and our model are due to densification differences, as memory per primitive is the same. FPS results are computed at test resolution, and results with a were not recomputed and use different high end GPUs.
We evaluate EVER on the 9 scenes introduced in Mip-NeRF 360 [3], the 4 scenes from DeepBlending [21] and Tanks&Temples [25], and the 4 scenes introduced in Zip-NeRF [4]. We use the same parameters across all scenes, except one: the size threshold for splitting and cloning is adjusted for large scenes to avoid total densification failure for both 3DGS and our method. For the alameda scene, we use the exposure trick introduced in [15] to allow evaluation on data sets with multiple exposures. We also re-tuned 3DGS to work slightly better than the tune done by Duckworth et al. [15] on the Zip-NeRF dataset.
We evaluate EVER on the 9 scenes introduced in Mip-NeRF 360 [3], the 4 scenes from DeepBlending [21] and Tanks&Temples [25], and the 4 scenes introduced in Zip-NeRF [4]. We use the same parameters across all scenes, except one: the size threshold for splitting and cloning is adjusted for large scenes to avoid total densification failure for both 3DGS and our method. For the alameda scene, we use the exposure trick introduced in [15] to allow evaluation on data sets with multiple exposures. We also re-tuned 3DGS to work slightly better than the tune done by Duckworth et al. [15] on the Zip-NeRF dataset.
There have been many follow-ups to 3DGS, each introducing different kinds of improvement, so we limit the scope of our comparison to best isolate our contribution. There are many recent approaches for densification [7, 43, 8, 50, 52], but we chose to change densification only in ways required to handle density-based parameterization. It is likely that recent progress on improved 3DGS heuristics may similarly improve the performance of our model. There are also many different ways to change rendering [18, 36, 5, 53, 22, 24], so we chose to only compare rendering methods that specifically address popping.
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption      Refer to caption   Refer to caption
3DGS StopThePop SMERF Ours GT
Figure 6: Visual comparison of our method on the Mip-NeRF [2] and Zip-NeRF [4] datasets. Rows 1 and 2 show regions where our results are sharper. Rows 3 and 4 show how proper blending of primitives helps with handling lighting on textureless surfaces. 3DGRT is omitted due to missing Zip-NeRF results.
6.1 Reconstruction Quality
We measure image quality through the standard image metrics PSNR, SSIM and LPIPS. We compare our method with 3DGS [23], and with Zip-NeRF [4], the current offline rendering state-of-the-art for these datasets. We also compare with StopThePop since it improves view-consistency for 3DGS, 3DGRT [32] because our work uses their ray tracing algorithm, and with SMERF [15], because even though it is a distillation based method that takes 33 hours to train on 16xA100 GPUs compared to 1-2 hours for our method, it is a comparable real-time 3D consistent volume rendering method.
Quantitatively, we outperform the other 3DGS-based methods across the board, and set a new state of the art in the Zip-NeRF dataset in terms of sharpness as measured by LPIPS and SSIM — even when compared to offline and distillation-based methods. Qualitatively, we notice two main factors: improved blending, and improved sharpness. The combination of more flexible primitive appearance, combined with color blending, allows our method to represent gradients better, which can be seen in Figure 4. This effect can also be seen when comparing how these methods model light falling on flat walls, as shown by the bottom two rows of Figure 6. While our method is able to reproduce the smooth image gradients in the shadowed areas, all of the Gaussian based methods struggle. The improved sharpness is more apparent on the larger and more difficult Zip-NeRF dataset, both in the metrics and in qualitative examples (see the second two rows of Figure 6).
6.2 Popping
StP Popping DeepBlending M360 Indoor M360 Outdoor Tanks & Temples
Metric
F
LIP1
F
LIP7
F
LIP1
F
LIP7
F
LIP1
F
LIP7
F
LIP1
F
LIP7
3DGS 0.0063 0.0122 0.0072 0.0149 0.0083 0.0154 0.0107 0.0315
StopThePop 0.0052 0.0055 0.0060 0.0073 0.0083 0.0115 0.0076 0.0114
Ours 0.0031 0.0000 0.0049 0.0000 0.0067 0.0000 0.0040 0.0000
Table 2: Experimental evidence that our method does not suffer popping, using the metric from StopThePop [36]. F LIP7 is considered a more reliable step size by the authors because the signal is stronger than the noise caused by the optical flow used to compute the metric.
To accompany our theoretical explanation of why our method does not suffer popping, we compare our method to StopThePop and 3DGS using the metric from StopThePop (StP) to measure the amount of popping, closely following the original author’s methodology across all the original scenes measured. EVER achieves a lower (better) score at step=1, and 0.0 error (lowest) with step=7, which StP called a more reliable step size, see Table. 2.
6.3 Performance
At 720p, we achieve average framerates of 36 FPS on the mip-NeRF 360 outdoor scenes, 66 FPS on the mip-NeRF 360 indoor scenes, and 30 FPS on the Zip-NeRF scenes on an NVIDIA RTX4090. The test set resolution is high on a couple of the Zip-NeRF and Mip-NeRF360 scenes, which causes the drop in FPS in Table 1. Training takes around 1-2 hours, also on an NVIDIA RTX4090. The majority of the time is spent tracing through the scene for rendering, which takes 20-71 ms, depending on the BVH. For back propagation, 16 ms are spent storing intersections, and 13-20 ms are spent loading the primitives and atomically adding up the gradients. The BVH is rebuilt every training iteration, which takes 2-20 ms.